A company that produces pleasure boats has decided to expand one of its lines. Current facilities are insufficient to handle the increased workload, so the company is considering three alternatives, A (new location), B (subcontract), and C (expand existing facilities).
Alternative A would involve substantial fixed costs but relatively low variable costs: fixed costs would be $290,000 per year, and variable costs would be $570 per boat. Subcontracting would involve a cost per boat of $2,520, and expansion would require an annual fixed cost of $59,000 and a variable cost of $1,110 per boat.

a. Find the range of output for each alternative that would yield the lowest total cost. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to the nearest whole number.)

A or more
B to
C to


b. Which alternative would yield the lowest total cost for an expected annual volume of 170 boats?

  • A

  • B

  • C
Answer #1

INDIFFERENCE POINT = (FIXED COST A - FIXED COST B) / (VARIABLE COST B - VARIABLE COST A)


FC A(NEW LOCATION) = 290000
VC A = 570
FC B(SUB CONTRACT) = 0
VC B = 2520
FC C(EXPAND) = 59000
VC C = 1110

INDIFFERENCE POINT(A TO B) = (290000 - 0) / (2520 - 570) = 149
INDIFFERENCE POINT(B TO C) = (0 - 59000) / (1110 - 2520) = 42

FOR A(NEW LOCATION), OPTIMAL QUANTITY = 149 OR MORE
B = 42 - 149
C = 0 - 42

B. ALTERNATIVE A, BECAUSE OF ITS LOWER VARIABLE COST PER UNIT, ABOVE THE CROSSOVER POINT OF 149 UNITS

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.